OpenStudy (kimberly101):

Use equations to solve the following questions Yolanda and Janelle play netball together and between them they have scored 103 goals, but Janelle has scored 15 more goals than Yolanda. How many goals has Yolanda scored?

OpenStudy (kropot72):

Let Y = Yolanda's score Then Janelle's score = Y + 15 Y + Y + 15 = 103 2Y + 15 = 103 Can you solve for Y ?

OpenStudy (kimberly101):

dont worry i got it the answer is 44

OpenStudy (kimberly101):

*the

OpenStudy (kimberly101):

thanks anyway

OpenStudy (kropot72):

Your answer is correct.

OpenStudy (kimberly101):

In a mathematics test Duy scored twice as many marks as Paul, but three marks less than Kate. The three students scored a total of 178 marks. How much did each one score?